OpenStudy (anonymous):

solve for h, S=2pir^2+2pir^1h

OpenStudy (anonymous):

wanna help me on this one too please?:) @gypsy1274

OpenStudy (anonymous):

If only I could....what is pir?

OpenStudy (anonymous):

\(S=2\pi r^2+2\pi r^1h\)

OpenStudy (anonymous):

yes thats it

OpenStudy (anonymous):

If the question read \(S=x+yh\) could you solve that for h?

OpenStudy (anonymous):

yes you would subtract x then divide by y to get \[h=\frac{ -x }{ y } \frac{ s }{ y }\]

OpenStudy (anonymous):

Almost. You would subtract x to get \(S-x=hy\) then divide by y to get \(\dfrac{S-x}{y}=h\). You could separate the fractions, but I seen no need for that.

OpenStudy (anonymous):

okay so now back to my question cause that one didnt involve all these exponents and pi. i know the technique but some im just so lost

OpenStudy (anonymous):

It is exactly the same thing. First subtract the term that does not include h.

OpenStudy (anonymous):

Then divide by everything but the h in the term that does have h. This one cannot be simplified any further.

OpenStudy (anonymous):

so it would be h=\[\frac{ s-2 \pi r ^{2} }{ 2 \pi r ^{1} }\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

wow i feel so dumb it all makes sense now thanks again :)